# Culturally Aware and Adapted NLP: A Taxonomy and a Survey of the State of the Art

### v1

@@ Abstract
The surge of interest in culturally aware and adapted Natural Language Processing (NLP) has inspired much recent research. However, the lack of common understanding of the concept of “culture” has made it difficult to evaluate progress in this emerging area. Drawing on prior research in NLP and related fields, we propose an extensive taxonomy of elements of culture that can provide a systematic framework for analyzing and understanding research progress. Using the taxonomy, we survey existing resources and models for culturally aware and adapted NLP, providing an overview of the state of the art and the research gaps that still need to be filled.
@@ 1 Introduction
Culture is rapidly becoming an important research topic in Natural Language Processing (NLP), with a significant recent surge in the number of published papers (Figure 1). Given the keen interest in this area and its importance for the safety and fairness of Large Language Models (LLMs), it is now important to consolidate existing research on culturally aware and adapted NLP to take stock of the progress made so far and to identify research gaps. However, this is challenged by the fact that there is no common understanding of the concept of “culture” in NLP.
@@
Prior work such as Hershcovich et al. (2022); Hovy and Yang (2021) laid the vital foundations for understanding ways in which language, culture and society interact in NLP. Hershcovich et al. (2022) proposed a simple taxonomy derived from the interaction between language and culture that captures some elements of culture (linguistic form and style, objectives and values, common ground, and aboutness). Recent papers have adopted a diversity of other definitions (using e.g., “proxies of culture”,  Adilazuarda et al. 2024), making it challenging to assess progress in this area.
Figure 1: Papers published on arXiv (cs.CL) that contain the keyword “culture”, aggregated every 3 months.
Figure 2: An overview of the taxonomy.
Although culture is arguably a complex concept, we can establish a common understanding by drawing on the definitions developed within the humanities and social sciences. In these fields Tylor (1871); Kroeber and Kluckhohn (1952); White (1959); UNESCO (1982); Matsumoto and Juang (1996); Blake (2000), most definitions of culture encompass people, groups of people and the interactions among both individuals and groups. [^1] Differences between cultures can exist within groups, such as due to languages; however, variations in language do not imply cultural differences in certain aspects and vice-versa (e.g., United Kingdom versus United States).
In anthropology, White (1959) provides valuable insights into the locus of culture, identifying three key dimensions: 1) “within human” (such as concepts, traditions, beliefs, social practices, etc.), 2) between “social interaction among human beings,” and 3) outside of human but “within the patterns of social interaction” (in materialized objects such as tools, arts). An examination of such work reveals that not only people, but also social interaction are critical components of culture.
In NLP, understanding of culture could benefit from the granularity needed to distinguish between more nuanced differences among cultural elements. Additionally, although social interactions are important, they have not been explicitly defined or categorized in prior work on cross-cultural NLP Hershcovich et al. (2022).
@@
We introduce a new extensive taxonomy of the elements of culture that addresses these problems. Expanding on the basic categories of prior work Hershcovich et al. (2022); Hovy and Yang (2021) and grounding the definition of culture in the relevant anthropology literature Tylor (1871); Kroeber and Kluckhohn (1952); White (1959), we introduce a fine-grained taxonomy that also covers social interactions. We then use this taxonomy to organize and analyze existing works in culturally aware and adapted NLP according to the elements of culture. Our survey of over a hundred publications in NLP provides an up-to-date view of cultural adaptation resources and models and identifies areas of progress as well as research gaps. We hope our taxonomy and analysis can enable and inspire future research in this emerging area.
@@ 2 The Taxonomy
In this section, we present our new taxonomy of culture. It differs from earlier NLP works Hershcovich et al. (2022); Hovy and Yang (2021) that aimed to define elements of culture in that (i) it is grounded in well-established elements of culture in anthropology literature Tylor (1871); Kroeber and Kluckhohn (1952); White (1959), (ii) consists of more fine-grained elements than in earlier work, and (iii) allows for a wider consideration of how social factors and variations in humans influence culture.
The taxonomy – shown in Figure 2 has two main branches: cultural elements3.1) within humans and outside of humans but within the social context White (1959) and sociocultural elements3.2) which affect and are affected by cultural elements through social interactions and communication among humans. [^2] Note that all elements of culture may interact and/or affect each other depending on context. Our taxonomy abstracts away from this contextual variation.
@@
Cultural Elements. Cultural elements are based on well-established definitions of culture Tylor (1871); Kroeber and Kluckhohn (1952); White (1959):
Concepts: basic units of meaning underlying objects, ideas, or beliefs Jackendoff (2012).
@@
Knowledge: information that can be acquired through education or practical experience.
Values: beliefs, desirable end states or behaviours ranked by relative importance that can guide evaluations of things Schwartz (1992).
@@
Norms and Morals: set of rules or principles that govern people’s behaviour and everyday reasoning Cialdini et al. (1991); Hechter and Opp (2001); Bicchieri et al. (2018); Gert and Gert (2002).
Linguistic Form: “how” to construct an utterance (Hershcovich et al., 2022, e.g., dialects).
Artifacts: “materialized” items as the productions of human culture, they can be forms of art, tools, machines White (1959) etc.
@@
Sociocultural Elements. “Sociocultural” elements refer to social factors within the scope of NLP that influence and are influenced by previously discussed elements. Leveraging the work of Hovy and Yang (2021), we identify relevant elements:
@@
Relationship: connection between two or more individuals or groups (e.g., father-son, colleagues).
@@
Context: the “containers” of communications Yang (2019), which can be linguistic such as surrounding sentences or extra-linguistic Hovy and Yang (2021) including social settings (e.g., at a wedding), non-verbal cues (e.g., gesture), or historical contexts (e.g., colonization).
@@
Communicative Goals: the intention behind language use (e.g., requests, apologies).
@@
Demographics: the characteristics of people, such as economic income, education level, nationality, location, political view, family status, etc.
@@ 3 Elements of Culture in Current NLP (Resources) Literature
In this section, we survey and categorize NLP resources published in leading *CL venues since 2022 (details in Appendix A) according to our new taxonomy (§2). For an overview of papers published in each cultural element class of the taxonomy, see Table 1 in the Appendix.
We observe that in resources, culture can be captured in 1) the data itself, or 2) in the labels (e.g., multi-culturally annotated).
3.1 Cultural Elements
@@ 3.1.1 Concepts
Recent NLP works have studied the grounding of time expressions to specific hours across different cultures, as detailed in Shwartz (2022, TimeExpressions, 27 languages). Cao et al. (2024b, CulturalRecipes, 2 languages) study culinary and food-related concepts within the context of recipe adaptations. Majewska et al. (2023); Hu et al. (2023) (4 languages each) culturally adapt names and locations for multilingual Task-Oriented Dialogue (TOD) datasets to enable cultural adaptation of TOD systems. Nevertheless, both training and evaluation datasets are still lacking in this area for diverse cultures, languages, and concept categories (e.g., rituals, aesthetics, kinship, spatial relations).
NLP research has also investigated the usage of concepts spanning multiple categories across cultures, e.g., through metaphors and other figurative expressions Kabra et al. (2023) or traditional proverbs and sayings Liu et al. (2023b).
In vision and language (VL) settings, culturally significant or unique concepts have been incorporated into reasoning and captioning tasks such as in MaRVL Liu et al. (2021), GD-VCR Yin et al. (2021), and XM3600 Thapliyal et al. (2022). Additionally, these concepts have been employed to evaluate multimodal content adaptations Khanuja et al. (2024) or to generate multicultural images Liu et al. (2023e); Ventura et al. (2023) with text-to-image models. These datasets are limited in size (e.g., due to the high cost of annotations) and most are available for evaluation only.
@@ 3.1.2 Knowledge
Cultural knowledge can be factual or common sense. [^5] Common sense and norms are sometimes used interchangeably in NLP. Norms are acceptable behavioural patterns of a group (§2), which we will discuss in §3.1.4. What weather phenomena can be expected if a rapidly rotating tropical storm forms off the coast of our country? (It’s likely called a hurricane in the US and a typhoon in Korea.) Is tofu pudding sweet or savoury by default? (In China, tofu pudding is typically sweet in the south and savoury in the north.)
@@
mLAMA Kassner et al. (2021), GeoMLAMA Yin et al. (2022), DLAMA Keleg and Magdy (2023), and FMLAMA Zhou et al. (2024) probe the diversity of knowledge in LMs.
Recently, MMLU Hendrycks et al. (2021) style QA benchmarks have greatly aided LLM development, inspiring many cultural variants of the MMLU suite. These suites include ArabicMMLU Koto et al. (2024a), CMMLU Li et al. (2023b), IndoMMLU Koto et al. (2023), SeaEval Wang et al. (2023a), KMMLU Son et al. (2024), CLIcK Kim et al. (2024), just to name a few. The cultural variants of MMLU typically cover some aspects of food, history, literature, geographical knowledge of countries, etc. in respective languages. However, MMLU-style benchmarks are typically constructed using exams and textbooks and cannot combine with other elements in culture. In contrast, common sense datasets (COPAL-ID, Wibowo et al. 2023 or IndoCulture, Koto et al. 2024b) can consider dialects and/or geographical diversity.
Lastly, CANDLE Nguyen et al. (2023), CultureAtlas Fung et al. (2024), MANGO Nguyen et al. (2024) are recent cultural knowledge bases created from either Wikipedia articles or using model distillation. They can be integrated with models to enhance models’ cultural awareness. However, diverse sources (i.e., other than Wikipedia, such as from local encyclopedia) should be considered to improve the coverage of knowledge bases.
@@ 3.1.3 Values
Many recent studies on evaluation (Johnson et al., 2022; Ramezani and Xu, 2023; Cao et al., 2023; Wang et al., 2023b; Durmus et al., 2023; Santurkar et al., 2023; Masoud et al., 2023; Havaldar et al., 2023, inter alia) show that LLMs align better with values of WEIRD (Western, Educated, Industrialized, Rich and Democratic, Henrich et al. 2010) people, raising concerns about the fairness and safety of LLM for others. Here, Pew Global Attitudes Survey (PEW) [^6] https://www.pewresearch.org/, the World Values Survey (WVS) [^7] https://www.worldvaluessurvey.org/ and the Hofstede Cultural Dimensions Hofstede (1984, 2011) are commonly used for evaluation, along with regional variants like the European Values Survey (EVS, EVS 2011). However, the questions of how to improve the model’s value alignment with diverse cultures, what resources to collect and whom to collect from remain unsolved Kirk et al. (2024).
@@
Biases. In contrast to cultural values, biases have been long-studied in NLP, such as gender bias in machine translation (Stanovsky et al., 2019; Savoldi et al., 2021; Campolungo et al., 2022; Sandoval et al., 2023; Attanasio et al., 2023, inter alia) or bias towards particular social groups. Cultures can have different biases toward the same target groups or have unique biases that are more pronounced in certain cultures (e.g., castes, unnatural beauty standards).
To enable evaluations of cross-cultural variations in biases and develop transferable de-biasing methods, recent work has created culturally aware targets and attribute word sets. WEATHub Mukherjee et al. (2023) and CA-WEAT España-Bonet and Barrón-Cedeño (2022) are multilingual culturally aware extensions based on the standard WEAT Caliskan et al. (2017) categories. Alternatively, there are other cultural variants such as BIBED (identity-bias in Bengali, Das et al. 2023), CAMeL (Arabic and Western cultural entities, Naous et al. 2023), SeeGULL (regional and national identity-bias, Bhutani et al. 2024), SPICE (stereotypes in India in English, Dev et al. 2023), CHBias (gender/orientation/age/appearance biases, Chinese, Zhao et al. 2023) among others.
In addition to using word lists, bias evaluation datasets may also consist of sentence pairs (such as French CrowS-Pair, Névéol et al. 2022; WinoQueer, Felkner et al. 2023), biased sentence and context (KosBi, Lee et al. 2023a), conversational (CDIAL-BIAS, Zhou et al. 2022) or question answering (FORK: food-related customs, Palta and Rudinger 2023; SODAPOP, An et al. 2023; KoBBQ, Jin et al. 2024).
@@
In general, there is considerable progress in cultural biases compared to other sub-areas. Many recent surveys on general biases are also available covering critical areas such as evaluations or de-biasing methods Meade et al. (2022); Dev et al. (2022); Sun et al. (2019); Delobelle et al. (2022), we refer readers to them for further details.
@@
Hate. The perception of hatefulness in text varies across cultures, languages and dialects, as recently shown for hate speech classification by Zhou et al. (2023a, b); Lee et al. (2023c); Lwowski et al. (2022); Arango Monnar et al. (2022), among others. Such model disparities may be due to the labelling process, where annotations from diverse cultural groups could serve as a mitigation strategy. CREHate Lee et al. (2023b) provides hate speech annotations from various English-speaking countries, all applied to the same English dataset. This example is unique in recent research as it explores variations in perceptions within the same language, highlighting the need for further research.
@@
Other Perceptions. The perception of politeness, aesthetic appeal or emotions can also vary across cultures House and Kasper (1981); Ringel et al. (2019); Masuda et al. (2008); Mesquita et al. (1997). For example, whether a piece of text is deemed humorous or ironic is culturally dependent. Frenda et al. (2023) tries to address this with a cross-cultural annotated irony corpora (EPIC) by people from five English-speaking countries. Similarly, visual elements in arts can elicit different emotions in different cultural groups. ArtELingo Mohamed et al. (2022) introduces a benchmark with Chinese, Arabic and Spanish captions and emotion labels for artworks, aimed at evaluating models’ cultural-transfer performance. This research area is significantly limited.

### v2

@@ Abstract
The surge of interest in culture in NLP has inspired much recent research, but a shared understanding of “culture” remains unclear, making it difficult to evaluate progress in this emerging area. Drawing on prior research in NLP and related fields, we propose a fine-grained taxonomy of elements in culture that can provide a systematic framework for analyzing and understanding research progress. Using the taxonomy, we survey existing resources and methods for culturally aware and adapted NLP, providing an overview of the state of the art and the research gaps that still need to be filled.
@@ 1 Introduction
Culture is rapidly becoming an important research topic in Natural Language Processing (NLP), with a significant recent surge in the number of published papers (Figure 1). Current NLP systems, especially Large Language Models (LLMs) often lack fairness and diversity in cultural awareness, which leads to biased performance that disproportionately favours certain groups, and causes harm to others Sambasivan et al. (2021); Johnson et al. (2022); Cao et al. (2023); Hofmann et al. (2024b). To build technology that is equitable, inclusive and accessible, the NLP community must actively take the initiative, enhancing LLMs’ cultural awareness and adaptability. Given the keen interest in this area and its importance for the safety and fairness of LLMs, it is now important to consolidate existing research on culturally aware and adapted NLP to take stock of the progress made so far and to identify research gaps. However, this is challenged by the lack of a common understanding of the concept of “culture” in NLP.
Figure 1: Papers title and abstract containing "culture" or "cultural" published in the main and findings of AACL/EACL/NAACL/ACL/EMNLP and TACL within 5 years, with normalized percentages based on the total number of papers at included venues to date.
@@
Prior work in NLP such as Hershcovich et al. (2022) laid the vital foundations for understanding how language, culture and society interact. Hershcovich et al. (2022) proposed a simple taxonomy derived from the interaction between language and culture that captures broad elements of culture (linguistic form and style, objectives and values, common ground, and aboutness). More recently, Adilazuarda et al. (2024) adopted “proxies of culture” (semantic or demographic proxies). While neither provides a shared understanding of culture, perhaps unsurprisingly, language is an essential component of culture in NLP.
Figure 2: An overview of the taxonomy with examples of subcategories of future possible expansions. The elements in culture are organized into three different branches: ideational, linguistic, and social. The ideational branch (§3.1 encompasses the non-material aspects of culture that constitute a way of life. The linguistic branch (§3.2) focuses on cultural variations in language and linguistic forms, bridging the ideational and social elements of culture. The social branch (§3.3) covers key factors in social interaction and communication.
A shared understanding of culture in NLP could benefit from examining definitions developed in anthropology and social sciences. [^1] They have been thinking about culture for over a century! In these fields (Tylor, 1871; Kroeber and Kluckhohn, 1952; White, 1959; UNESCO, 1982; Matsumoto and Juang, 1996; Blake, 2000; Geertz, 1973; Goffman, 2023), most definitions of culture involve people, groups, and the interactions within and between individuals and groups.
Murdock (1940) describes culture as “ideational” [^2] The values, beliefs, norms, and ideas that comprise a way of life Murdock (1940); Briggle and Mitcham (2012b). and “social”. White (1959) describes the locus of culture as: 1)within human” (e.g., concepts, beliefs, i.e., ideational), 2) between “social interaction among human beings”, and 3) outside of human but “within the patterns of social interaction”. An examination of such work reveals that social interactions are critical components of culture, in addition to ideational elements. Notably, this social aspect is underrepresented in previous cross-cultural NLP research Hershcovich et al. (2022); Adilazuarda et al. (2024). [^3] There is also prior work focusing on social NLP Hovy and Yang (2021) which is related. However, culture is not the central theme of the work.
Hence, we combine the emphasis on language in prior NLP work, while integrating the significance of social and ideational elements that shape culture. This leads us to a working definition of culture used to taxonomize recent work in NLP in this survey: Culture encompasses the collective ideas, shared language, and social practices that emerge from and evolve through human social interactions within a society.
@@
Grounded in this working definition, we introduce a new fine-grained taxonomy of culture by expanding on the basic categories of prior work Hovy and Yang (2021); Hershcovich et al. (2022) to address above-mentioned issues. We then use this taxonomy to organize existing works in culturally aware and adapted NLP, and identify research gaps. Our survey of 127 publications in leading *CL venues (see selection method in Appendix A) provides an up-to-date view of cultural adaptation resources and models and identifies areas of progress as well as new research opportunities. We hope our taxonomy and analysis enable and inspire further research in this important emerging area.
@@ 2 The Taxonomy
In this section, we present our new taxonomy of culture. Unlike previous NLP studies Hershcovich et al. (2022); Hovy and Yang (2021) that sought to define cultural elements, this taxonomy (i) is grounded in well-established elements of culture in anthropology and social sciences Tylor (1871); Kroeber and Kluckhohn (1952); White (1959), (ii) consists of more fine-grained elements than in earlier work, and (iii) allows for a wider consideration of how social factors and variations in humans influence culture.
Figure 2 presents a taxonomy of cultural elements derived from our working definition in §1, organized into three main branches: ideational, linguistic, and social. [^4] Icons in Figure 2 are created with the assistance of DALL-E. The ideational branch (§3.1; Murdock 1940; Briggle and Mitcham 2012a) encompasses the non-material aspects of culture that constitute a way of life, such as values or knowledge. The linguistic branch (§3.2) focuses on cultural variations in language and linguistic forms, bridging the ideational and social elements of culture. The social branch (§3.3) covers key factors in social interaction and communication, such as relationships or communicative goals. [^5] All cultural elements can interact and influence each other based on context and can be divided into finer groups. Similar to prior work, our taxonomy abstracts away from these contextual variations. Here, we define each element based on existing research and relating to example tasks in the NLP context. We then provide details and examples from the current literature in §3.1, §3.2 and §3.3.
@@
Ideational elements are based on well-established discussions of culture Tylor (1871); Kroeber and Kluckhohn (1952); White (1959):
Concepts: basic units of meaning that structure and facilitate thought, bridging sensory experience Jackendoff (1989, 2012), e.g., cuisines (such as schnitzel, ratatouille) or holidays (such as Diwali, Nowruz). Related NLP task examples: question answering, dialogue generation.
@@
Knowledge: information that can be acquired through education or practical experience, e.g., local agricultural knowledge. Related NLP task examples: dialogue generation, reasoning.
Values: beliefs, desirable end states or behaviours ranked by importance that can guide evaluations of things Schwartz (1992). Unlike norms and morals, values do not inherently involve ethical judgment, e.g., beauty standards, or perception of hate speech. Related NLP task examples: content moderation, debiasing.
@@
Norms and Morals: set of rules or principles that govern people’s behaviour and everyday reasoning Cialdini et al. (1991); Bicchieri et al. (2018); Hechter and Opp (2001); Gert and Gert (2025), e.g., filial obedience attitude. Related NLP task examples: reasoning, safety alignment.
Artifacts: items that are products of human culture, such as art, poetry White (1959) etc. This is ideational in our taxonomy since we do not work on physical buildings or tools in NLP. Related NLP task examples: machine translation of long-form literature, emotion arc analysis of movies, memes classification.
Linguistic elements relate to language variations in the cultural context, based on existing discussions Wardhaugh and Fuller (2021):
Dialects: includes variations of languages in a systematic way (Fromkin et al. 1998; Trudgill 2000; Wardhaugh and Fuller 2021; such as dialects continuum, regionalects, sociolects etc.), e.g., African American (Vernacular) English (AAE/AAVE). Related NLP task examples: machine translation, debiasing.
Styles, Registers, Genres: includes elements such as formality, variations of language in situation and communicative forms Wardhaugh and Fuller (2021), e.g., formality in text, slang, or specific genres like news, folk tales. Related NLP task examples: style transfer, creative writing generation.
@@
Social elements focus on social interactions and communication among humans within the scope of NLP. Leveraging the work of  (Hovy and Yang, 2021), we identify relevant elements:
@@
Relationship: connection between two or more individuals or groups, e.g., father-son, colleagues. Related NLP task examples: creative writing generation, dialogue generation.
@@
Context: the “containers” of communications Yang (2019), which can be linguistic such as surrounding sentences or extra-linguistic Hovy and Yang (2021) including social settings (e.g., at a wedding), non-verbal cues (e.g., gesture), or historical contexts (e.g., colonization). Related NLP task examples: coreference resolution, pragmatic inference.
@@
Communicative Goals: the intention behind language use, e.g., requests, apologies, persuasion. Related NLP task examples: intent classification, emotion classification, human-AI collaboration.
@@
Demographics: the characteristics of people, e.g., economic income, education level, nationality, location, political view, family status, etc. Related NLP task examples: content moderation, personalization.
@@ 3 Elements of Culture in Current NLP (Resources) Literature
In this section, we survey and categorize NLP resources. Table 1 shows an overview of papers organized according to the taxonomy.
We observe that in resources, culture can be captured in 1) the data itself, or 2) in the labels (e.g., multi-culturally annotated). Further, while cultural differences are evident in linguistic and social elements, most current work relies on standard language or country boundaries, leaving these elements understudied.
Element Papers
Concepts Shwartz 2022; Majewska et al. 2023; Hu et al. 2023; Kabra et al. 2023 Liu et al. 2024a; Cao et al. 2024b; Jiang and Joshi 2024; Hu et al. 2024 Vision-Language: Liu et al. 2021; Yin et al. 2021; Thapliyal et al. 2022 Khanuja et al. 2024; Li et al. 2024c; Bhatia et al. 2024; Nayak et al. 2024
[…]
Dialects Ziems et al. 2022a, 2023b; Le and Luu 2023; Plüss et al. 2023; Kuparinen et al. 2023 Elmadany et al. 2023; Deas et al. 2023; Khondaker et al. 2023; Faisal et al. 2024
[…]
Relationship Li et al. 2023b; Jurgens et al. 2023; Shaikh et al. 2023; Ziems et al. 2023a; Zhan et al. 2024
[…]
Table 1: Recent resource work considered in §3 by elements (selection method in Appendix A). The three blocks (divided by double lines) correspond to ideational, linguistic, and social elements respectively.
3.1 Ideational Elements
@@ 3.1.1 Concepts
Recent NLP research has explored grounding time expressions across cultures (Shwartz, 2022) and culinary concepts in recipe adaptations (Cao et al., 2024b). Additionally, studies have examined how various cultures use concepts across categories, such as through metaphors Kabra et al. (2023) or traditional proverbs and sayings Liu et al. (2024a). In vision and language (VL) settings, culturally unique concepts have been integrated into reasoning and captioning tasks (Liu et al., 2021; Yin et al., 2021; Thapliyal et al., 2022; Li et al., 2024c) or assess multimodal content adaptations Khanuja et al. (2024) and generation of text-to-image models Liu et al. (2023c).
These datasets are often small due to high annotation costs, and most are only available for evaluation. Training and evaluation datasets still lack diversity across cultures, languages, and concept categories (e.g., rituals, aesthetics, spatial relations).
@@ 3.1.2 Knowledge
Cultural knowledge can be factual or common sense. [^8] Common sense and norms are sometimes used interchangeably in NLP. Norms are acceptable behavioural patterns of a group (§2), which we will discuss in §3.1.4. What weather phenomena can be expected if a rapidly rotating tropical storm forms off the coast of our country? (It’s likely called a hurricane if one is in US, a typhoon if one is in Korea.) Is tofu pudding sweet or salty by default? (In China, it’s typically sweet in the south but salty in the north.)
@@
Assessing language models’ knowledge has long been important, with early studies examining this across languages predating LLMs Kassner et al. (2021); Yin et al. (2022); Keleg and Magdy (2023); Zhou et al. (2024). Recently, MMLU-style Hendrycks et al. (2021) MCQA benchmarks have advanced LLM development and inspired cultural variants (e.g., Li et al. 2024b, details in Table 1) covering aspects like food, history, and geography in respective languages. However, MMLU-style benchmarks, often based on standard exams and textbooks, lack integration with broader cultural elements. In contrast, other common sense knowledge datasets (e.g., Wibowo et al. 2024, Koto et al. 2024b) can incorporate other elements under “linguistic” (e.g., in dialects) or “social” (such as from diverse demographics with geographic regions) branches.
Finally, integrating knowledge bases (KB) with models enhances cultural awareness Bhatia and Shwartz (2023) and supports culturally relevant synthetic data generation Kim et al. (2023). Despite recent efforts in creating cultural KB from other venues Nguyen et al. (2023); Fung et al. (2024); Nguyen et al. (2024), *CL community examples remain limited.
@@ 3.1.3 Values
Many recent studies on evaluation (Johnson et al., 2022; Ramezani and Xu, 2023; Cao et al., 2023; Durmus et al., 2024; Santurkar et al., 2023; Masoud et al., 2025; Havaldar et al., 2023b; Wang et al., 2024c, inter alia) show that LLMs align better with values of WEIRD (Western, Educated, Industrialized, Rich and Democratic, Henrich et al. 2010) people, raising concerns about the fairness and safety of LLMs for others. Here, Pew Global Attitudes Survey (PEW) [^9] https://www.pewresearch.org/, the World Values Survey (WVS) [^10] https://www.worldvaluessurvey.org/ and the Hofstede Cultural Dimensions Hofstede (1984, 2011) are commonly used for evaluation, along with regional variants like the European Values Survey (EVS, EVS 2011). However, the questions of how to improve the model’s value alignment with diverse cultures, what resources to collect and whom to collect from remain unsolved Kirk et al. (2024).
@@
Biases. In contrast to general cultural values, biases have been long-studied in NLP, such as gender bias in machine translation (Stanovsky et al., 2019; Savoldi et al., 2021; Campolungo et al., 2022; Sandoval et al., 2023; Attanasio et al., 2023, inter alia) or bias towards particular social groups. Differences in value “ranking” lead cultures to exhibit distinct biases toward the same groups or unique biases specific to certain cultures (e.g., caste systems, unnatural beauty standards). These variations are central to the study of cultural biases and are the focus of our work.
To enable evaluations of cross-cultural variations in biases and develop transferable de-biasing methods, recent work has created varieties of culturally aware datasets to aid evaluations, including targets and attribute word sets, sentence pairs, conversational and QA data (see Table 1 for the papers).
@@
Overall, this area shows notable progress compared to other sub-areas. Recent surveys on general biases cover key topics like evaluation and de-biasing methods Sun et al. (2019); Meade et al. (2022); Dev et al. (2022); Delobelle et al. (2022), which we refer readers to them for further details.
@@
Hate. Like biases, perceptions of hatefulness in the text also vary across cultures, as shown in recent research on hate speech classification (Sap et al., 2022; Zhou et al., 2023a, b; Lee et al., 2023b; Lwowski et al., 2022; Arango Monnar et al., 2022). Such model disparities may be due to the data source (i.e., using machine translations, not native text) or the labelling process. The first issue can be addressed by diversifying data sources, incorporating authentic local data (Shekhar et al., 2022; Jeong et al., 2022). The second issue can be mitigated by creating annotations from diverse cultural groups. Recently, CREHate Lee et al. (2024b) investigates variations in hate speech perceptions within the same language, highlighting the need for further research.
@@
Other Perceptions. The perception of politeness, aesthetic appeal or emotions can also vary across cultures House and Kasper (1981); Mesquita et al. (1997); Masuda et al. (2008); Ringel et al. (2019); Abdelkadir et al. (2024). For example, whether a piece of text is deemed humorous or ironic is culturally dependent. Frenda et al. (2023); Casola et al. (2024) try to address this with cross-cultural annotated (multilingual) irony corpora. Similarly, visual elements in arts can elicit different emotions in different cultural groups. ArtELingo Mohamed et al. (2022, 2024) provides benchmarks with multilingual captions and emotion labels for artworks to evaluate models’ cultural-transfer performance. This research area is significantly limited.